OpenStudy (anonymous):

Use the limit definition to prove the limit lim x->3(x^2 -5x +1)=-5

OpenStudy (anonymous):

what definition are you using ?

OpenStudy (anonymous):

are you using \(\epsilon, \delta\) business?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

k then the first part is a bunch of algebra

OpenStudy (anonymous):

we want to show that given any \(\epsilon>0\) there is some \(\delta\) (which we write in terms of \(\epsilon\) so that if \(|x-3|<\dela\) we know that \(|x^2-5x+1+5|<\epsilon\)

OpenStudy (anonymous):

typo there, i meant if \(|x-3|<\delta\) ...

OpenStudy (anonymous):

can anyone help me plz too

OpenStudy (anonymous):

okay so i get the first part but how do you choose

OpenStudy (anonymous):

ughh...sorry still trying to get used to this

OpenStudy (anonymous):

how do i choose delta

OpenStudy (anonymous):

ok so are you good to this point? wanting to make \(|x^2-5x+1+5|<\epsilon\) ?

OpenStudy (anonymous):

yeah, and then i get \[\left| x-2 \right|\left| x-3 \right|\]

OpenStudy (anonymous):

oh then you are home free almost

OpenStudy (anonymous):

you get control over the \(|x-3|\) part, i.e. you can make it as small as you like

OpenStudy (anonymous):

now there is one little gimmick, what do to about the \(|x-2|\)

OpenStudy (anonymous):

you have many choices, but since you get to say how close \(x\) is to \(3\), i.e. how small \(|x-3|\) is, you can say it is certainly less than \(1\) if it is less than \(1\) that makes \(2<x<4\) in which case \(|x-2|\) is largest if \(x=4\) and \(|4-2|=2\) so at the very most \(|x-2||x-3|<2|x-3|\) if we assume \(|x-3|<1\)

OpenStudy (anonymous):

and since if \(\delta<1\) we have \(|x-2||x-3|<2\delta\) all we have to do to insure it is less than \(\epsilon\) is to make \(|x-3|<\frac{\epsilon}{2}\)

OpenStudy (anonymous):

does taking 1 work everytime, or just in this case?

OpenStudy (anonymous):

since you have total control over \(|x-3|\) you could have picked \(|x-3|<5\) if you like, really makes no difference, i just picked \(1\) because it was simplest to compute with and in any case we are considering small values of \(\epsilon\) not large ones

OpenStudy (anonymous):

thank you soo much..this really helped..

OpenStudy (anonymous):

you probably want to say if \(\delta<\text{min}(1,\frac{\epsilon}{2})\) then \(|x-2||x-3|<\epsilon\)

OpenStudy (anonymous):

and if your teacher is a stickler for details you have to run the algebra backwards for the complete proof, but it is really exactly that, writing the steps in reverse order yw

OpenStudy (anonymous):

My book hardly has any answers, so even if I am doing something right I'm not really sure because I have no way to check.. My prof's alright..but I'm trying to work on my proofs still..

OpenStudy (anonymous):

it is a whole new world, these proofs but that is actually all of math, at least all of real math

OpenStudy (anonymous):

Yeah this is the first time I'm doing proofs..don't like it at all!